OpenStudy (anonymous):

hydrogen peroxide can decompose to water and oxygen by the following reaction 2H2O2 (l) ------ 2H2O(l) + O2(g) enthalpy=-196kj calculate the value of q when 5.00g of h20(l) decomposes at constant pressure.

OpenStudy (anonymous):

1 mole of H2O2=34.0148 g 5.00g of H2O2=5/34.0148=0.146995 moles decomposition of 2moles produce -196kj therefore q produced by 0.146995 moles of H2O2=(-196/2)*0.146995=-14.40551Kj i think this is the way , i hve not been through this calculation for a long period~9-10years sorry if there is a mistake...........

OpenStudy (anonymous):

the answer is correct...thanx for the explanation